I have 80% solution (vinegar) and need 7% solution (Vinegar).
How much water do I have to add to get the result?

Found 2 solutions by richwmiller, amalm06:
Answer by richwmiller(17219)

Depends how much you have (x) or how much you want to have (x+y).
.80x=.07(x+y)
or from standpoint of water
.20x+1.0y=.93(x+y)

Answer by amalm06(224)

(0.80)(x)+0=(0.7)(x+w), where x is the amount of vinegar and w is the quantity of water
0.8x=0.7x+0.7w
0.1x=0.7w
w=(7/10)x (Answer)